OpenStudy (anonymous):

Solve for x: 4 − (x + 2) < −3(x + 4) x < −7 x > −7 x < −9 x > −9

OpenStudy (anonymous):

4 − (x + 2) < −3(x + 4) 4 − (x + 2) = −3(x + 4) 4-x-2=-3x-12 2=x-3x-12 14=-2x x=-7 replace x with -7 on both sides. 4 − (-7 + 2) < −3(-7+ 4) 5<33 so the answer would be x<-7